OpenStudy (anonymous):

2 1/3+3 1/2 a.5 1/5 b.5 2/5 c.4 9/10 d.4 1/5

OpenStudy (anonymous):

convert \[2\frac{ 1 }{ 3 } + 3\frac{ 1 }{ 2 }\] to improper fractions. It would become \[\frac{ 7 }{ 3 }+\frac{ 7 }{ 2 }\] Then Find the least common denominator, which is 6, and make both fractions' denominators 6.\[[\frac{ 7 }{ 3 }\times \frac{ 2 }{ 2 }]+[\frac{ 7 }{ 2 }\times \frac{ 3 }{ 3 }]= \frac{ 14 }{ 6 }+\frac{ 21 }{ 6 }=\frac{ 35 }{ 6 }\]

OpenStudy (anonymous):

35/6

OpenStudy (anonymous):

thank u :)